OpenStudy (anonymous):

the length of a rectangle is 16 more tahn twice its width. The area is 40 inches. Find the demensions of the rectangle

OpenStudy (bibby):

A = lw w l = 16 + 2 times w

OpenStudy (bibby):

You can call w, x. So we have: w = x l = 16 + 2x A = x(16+2x) = 40

OpenStudy (anonymous):

Thanks Bibby

OpenStudy (bibby):

No problem.